# Visual collective behaviors on spherical robots

### v2

@@ Abstract
The implementation of collective motion, traditionally, disregard the limited sensing capabilities of an individual, to instead assuming an omniscient perception of the environment. This study implements a visual flocking model in a “robot-in-the-loop” approach to reproduce these behaviors with a flock composed of 10 independent spherical robots. The model achieves robotic collective motion by only using panoramic visual information of each robot, such as retinal position, optical size and optic flow of the neighboring robots. We introduce a virtual anchor to confine the collective robotic movements so to avoid wall interactions. For the first time, a simple visual robot-in-the-loop approach succeed in reproducing several collective motion phases, in particular, swarming, and milling. Another milestone achieved with by this model is bridging the gap between simulation and physical experiments by demonstrating nearly identical behaviors in both environments with the same visual model. To conclude, we show that our minimal visual collective motion model is sufficient to recreate most collective behaviors on a robot-in-the-loop system that is scalable, behaves as numerical simulations predict and is easily comparable to traditional models.
@@ 1 Introduction
Biological collective motion has inspired efforts to understand, replicate, and ultimately utilize on artificial agents. Nearly 40 years ago, Reynolds et al.,[1] proposed that a set of simple, individual-level rules could explain collective motion. These rules include (i) attraction, the desire to remain close to other individuals; (ii) alignment, the tendency to move in the same direction as the group; and (iii) avoidance, the willingness to avoid collisions with others. Vicsek et al.,[2] later introduced a model based on statistical analysis, which averaged the velocities of individuals within a defined proximity of the individual and accounted for individual tendencies as stochastic noise. Couzin et al.,[3] expanded upon these ideas by combining the rules defined by Reynolds with the area of interaction defined by Vicsek. Authors propose that each rule have an area of interaction: the closest avoidance, the following alignment and the biggest one attraction. These 3 rules effectively reproduce in simulation three collective behaviors observed in nature, swarming (no orientational order), schooling (highly polarized movement), and milling (circular, swirling movement). Models that follow these guides are now known as “3A models”[4, 5, 6].
Most of available approaches to 3A-models rely on artificial tools to describe or enforce the rules, such as inter-individual distance[7, 8, 9], global velocity[10, 11], or absolute communication among individuals[12, 13]. However, these methods do not accurately reflect the way animals interact with their environment or others. To address this gap, researchers have explored biologically plausible (“bio-plausible”) ways of describe the environment[14, 15, 16, 17], with vision proving to be the most effective and widespread one[18, 19, 20, 21, 22]. Several vision-based models have been tested on robotic platforms[23, 24], but both of these either use vision merely as a selective mechanism[25, 26] or lack alignment with the fundamental principles of the 3A models[27].
Castro et al. (2024) [28] introduced a model that utilizes only visual information while maintaining the traditional 3A model structure. This model effectively replicates the phases of collective motion described by Couzin et al., by employing a visual mechanism known as optic flow. Optic flow, which captures the apparent angular motion of the environment due to the relative movement between an observer and its surroundings, has been validated as a crucial navigational tool in nature. Optic flow magnitude is used by insects during most flight maneuvers [29, 30] as well as in fish and birds [31, 32]. Furthermore, several robotic platforms have already incorporated small sensors dedicated to measuring optic flow, highlighting its potential as a key mechanism for both biological understanding and innovation[33, 34, 35, 36, 37].
@@
Building on these findings, this study proposes a “robot-in-the-loop” approach to collective motion that utilizes only visual information, implemented on a educational robotic platform. Unlike previous models, this approach does not rely on global information or complex communication networks. Instead, it introduces a visual anchor to account for environmental boundaries, thereby eliminating the need to model wall interactions explicitly. Experimental results and numerical simulations demonstrate that this method successfully reproduces several collective motion behaviors on a flock of 10 independently controlled robots.
@@
The article is organized as follows: Section 2.1 provides an overview of the challenges of a robotic implementation of 3A-models, with the example of the visual-based collective motion model proposed by Castro et al. (2024) [28] (Fig.1). Section 2.2 describes the robotic platform and the model modifications for implementing on the 10 independently controlled robots including the introduction of the visual anchor concept. Section 2.4 detail the implementation of a feature matching optic flow and visual collective motion rules. Section 2.5 describes the supplementation on original model and summarizes robotic setup, while Section 3 presents the results from simulations and robot-in-the-loop experiments. Finally, Section 4 concludes with a discussion of the results and directions for future research.
@@ 2 Methods
Figure 1: Model collective behaviors from optic flow and retinal cues [28]. a) View of the flock positions.b) Representation of the flock positions after recognition at time step $t$. Each $i$-particle is described by its XY-coordinates and unit heading vector $\mathbf{e}_{i}$, with the heading angle $\theta_{i}$. c) Point-of-view of particle 0 showing the retinal relative positions of each particle. The arrows represent the perceived optic flow, and the white arrow indicates the apparent velocity ($\mathbf{V}_{i,k}$). Here, we only see the azimuthal component of the apparent velocity $\mathbf{V}_{i,k}$ by the particle i=“0” of retinal object “k”. d) $\mathcal{V}(\phi)$ binary panorama representation of the field of view of particle 0. e) Apparent distance $\mathcal{R}_{i}(\phi)$ of particle “0”. f) Optic flow $\mathcal{O}_{i}(\phi)$ perceived by particle 1. g) Optic flow divergence $\mathcal{D}_{i}(\phi)$ perceived by particle 1.
@@ 2.1 Robotic Implementation challenges
Flocking models that are based on individual rules of attraction, avoidance and alignment, or better known as 3A-models, present several challenges when being implemented on a robotic platform. The biggest two are related to how individuals are modeled. The assumption of most models is that each individual is a self propelled particle, traditionally moving on a plane. Yet most robots are particles or restrict their movement only to a plane. This brings forth the first challenge, which robot kinematics should be chosen? It should balance following the assumptions of the model, sensing capabilities, computational capacity, size, among others. By exiting the numerical environment, the experimental environment are not infinite anymore, so the collective behavior should be confined which is a second challenge. Without this considerations flocking on robots would be as direct as a raw implementation of a established model, such as Castro et al., (2024) [28] on a the selected robotic platform (as seen on Fig. 1). The following sections will present our solution to both challenges.
@@ 2.2 Robotic setup
This section will section will present our robot setup to a robot-in-the-loop visual flocking model. We propose to use an underpowered, commercially available robotic platform that offloads the sensing capabilities to a base station but independently uses the information from its binary visual map to execute the model presented in the previous section.
@@
Figure 2: Robotic platform. a) Sphero Bolt®wireless educational robots. b) Example of the robots’ movement, c) Heading control by differential rotation on the two opposite inner wheels, d) Induced nose-up pitch during forward displacement.
@@
This robot-in-the-loop solution utilizes an existing educational robotic platform, the Sphero Bolt®(Fig. 2a). These are small, spherical robots that can be controlled in real-time via Bluetooth®and an educational app, where users can set a heading value, desired velocity, LED states, and read the onboard sensors’ values. Additionally, several open-source projects integrate these robots with different programming environments such as ROS, Python, or MATLAB®.
@@
Each robot is equipped with several sensors, including an Inertial Measurement Unit (IMU) comprising a 3-axis accelerometer, gyroscope, and magnetometer, and a light sensor. It also features four infrared (IR) proximity sensors on its sides, two addressable RGB LEDs (one in the front and one in the back), and an 8x8 RGB LED matrix on top.
@@
The Sphero’s movement is achieved via a self-balancing two-wheel enclosed system, with two non-actuated wheels maintaining constant contact with the enclosing sphere. This dual-motor design allows for stationary turning; however, while moving, the turning is achieved by differential motion (Fig. 2b-c). The locomotion generates a pitch displacement depending on the speed and the friction coefficient between the sphere’s encapsulation and the surface (Fig. 2d). This pitch is partially managed by the robot’s dual-MCU configuration: one MCU handles the communication protocol and high-level instructions, while the other manages motor control to ensure the robot achieves the desired heading and speed, along with low-level sensor readings.
@@
Figure 3: Experimental setup. a) Picture of the experimental setup, b) Example of the images taken by the camera (Daheng Imaging MER2-230-168U3C camera with 4mm c-mount lens), c) Whole robot body image segmentation, d) Individual light sources and heading recognition, e) Individual robot identification and virtualization.
@@
To address the lack of an onboard camera and the inability to embed one into the robots, we set up a bird’s-eye camera above the experimental area. This is an industrial RGB global shutter camera, the Daheng Imaging MER2-230-168U3C, with a frame rate of 168 FPS, a resolution of 1920x1200, and support for interchangeable c-mount lenses (Table 1). The selected lens is a 4mm, 1/8”, F2.0-12 non-distorting lens, which allows the camera to be mounted 2.3 m above the robots, covering a 2.7x2.7 m working area (approximately 38 times the length of the robots) (Fig. 3a).
@@
Table 1: Camera Settings.
Each robot is assigned an individual color pattern that enables identification of the robot’s front and back. To maintain light connectivity between the front and back, the back half of the LED matrix is set to 50% bright white light. This configuration provides 16 possible combinations for identifying each robot, with the potential for expansion using additional colors. The bird’s-eye camera allows the isolation of each light source on top of the robots, preventing light bleeding between robots when they are in close proximity (Fig. 3b). This is achieved by capturing a 1920x1200 RGB image and performing whole-robot segmentation. This segmentation treats the three light sources as a single entity, allowing calculation of the center’s XY-coordinates, which approximate the robot’s center. Knowing the radius of the robots in pixels allows for cropping of $N$ images from these coordinates, each isolating a single robot. The cropped images are further segmented to identify the front and back LEDs, as well as the LED matrix. The larger area generated by the two light sources enables front and back recognition (Fig. 3d). Additionally, by correlating the color from this segmentation, each robot can be uniquely identified. The final step is to compute the robot’s heading by tracing the vector between the center positions of the front and back LEDs. This process results in the identification, XY-coordinates, and heading of each robot (Fig. 3e).
The next step, after identifying each robot and determining their location and heading, is to reconstruct the binary panorama of each robot.
@@ 2.3 Anchoring the collective behavior
Figure 4: Binary panorama reconstruction. a) 2D virtual representation of the binary panorama, the red goal represents the desired location of the flock. b) Mixed reality point of view of the 0-particle with the omniscient red goal. c) Graph representation of all $\mathcal{V}(\phi,t)$. $\mathcal{V}(\phi,t)$ is a binary panorama representation of the field of view of particle i where the anchor is represented a different channel (red). This graph can be stored as a tuple list (rising and falling edges).
In the binary panorama reconstruction, we introduce a new component to the original model [28]: a visual anchor. The inclusion of an anchor bounds the flock to a place and prevent infinite expansion. The anchor is an augmented reality (AR) solution that enables each robot to see the desired center point (Fig. 4a). This anchor is omnisciently visible, regardless of any obstacles between the line of sight and the anchor location (Fig. 1b). It is a new visual channel that uses the same binary representation as the visual surroundings (Fig. 1c). The binary panorama is represented in Python as an ordered tuple list of every retinal object’s rising and falling edge, with the last position always being the goal tuple. This process is performed independently for each particle without sharing information between particles.
@@ 2.4 Robotic Visual Flocking Model
Avoidance effect ($\mathcal{A}$) is here introduced in the attraction term ($\omega_{\odot}$). Our previous study [28] did not include any avoidance because most collective motion models are based on the assumption that all particles have some extra room on the z-axis, allowing them to avoid collisions with additional depth. This assumption is primarily based on how fish flocking data is gathered, where the fish can swim above or below each other in a few centimeters of water [38]. However, for any robotic implementation, obstacle or collision avoidance is of critical importance. Fortunately, avoidance and attraction have similar characteristics, being proportional instantaneous responses to the optical size and location of a retinal object. The key difference is in the direction of the response: while attraction is positive, avoidance is negative. Therefore, avoidance is implemented by providing a new piecewise constant function ($\mathcal{A}$), where attraction is inverted in situations where a particle should avoid a retinal object. This new function depends on the width of the retinal object, specifically the difference between its rising and falling edges ($\Delta\phi$). This being a indirect measurement of the apparent distance ($\mathcal{R}(\phi,t)$). We choose $\mathcal{A}(\Delta\phi)=-5$ for $\Delta\phi<\pi/10$, which is equivalent to the retinal object being closer than three particle radii ($\mathcal{R}<3a$). In all other cases, $\mathcal{A}(\Delta\phi)=1$. The blind spot remains unchanged, giving more weight to the retinal object in front of the particle, regardless of this new function.
@@
$\frac{\mathbf{e}_{i}(t)\times\mathbf{V}_{ik}(t)}{\mathcal{R}_{i}(t)}=-\mathcal{D}_{i}(\phi,t)\sin(\phi,t)+\overline{\mathcal{O}}_{i}(\phi,t)\cos(\phi,t).$ (1f)
Lastly, the robotic platform introduces several critical processing factors – Bluetooth, image processing, internal Spheros control command, … – that affect the overall dynamical response of the robots and its collective behavior as a whole (see Supplementary Information for more details). These extra factors results in a time delay $\tau$ with a median of 600ms. The overall functional diagram of the collective motion model is presented in Fig. 5. This model is executed independently and simultaneously for all robots; it does not involve communication between robots, sharing information, and only uses visual data. The same block could be utilized on a robotic platform using LiDAR, onboard cameras, touch sensors, ultrasound, or any other sensor that provides a binary visual panorama as output. All the information used by this model is self-referred and self-contained; there is no global reference frame, coordination, or knowledge of the flock or the environment.

### v3

@@ Abstract
The implementation of collective motion, traditionally, disregard the limited sensing capabilities of an individual, to instead assuming an omniscient perception of the environment. This study implements a visual flocking model in a “robot-in-the-loop” approach to reproduce these behaviors with a flock composed of 10 independent spherical robots. The model achieves robotic collective motion by only using panoramic visual information of each robot, such as retinal position, optical size and optic flow of the neighboring robots. We introduce a virtual anchor to confine the collective robotic movements so to avoid wall interactions. For the first time, a simple visual robot-in-the-loop approach succeed in reproducing several collective motion phases, in particular, swarming, and milling. Another milestone achieved with by this model is bridging the gap between simulation and physical experiments by demonstrating nearly identical behaviors in both environments with the same visual model. To conclude, we show that our minimal visual collective motion model is sufficient to recreate most collective behaviors on a robot-in-the-loop system that be implemented using several individuals, behaves as numerical simulations predict and is easily comparable to traditional models.
@@ 1 Introduction
1.1 Related Work
Biological collective motion has inspired extensive efforts to understand, replicate, and utilize it in artificial systems. Nearly four decades ago, Reynolds et al. [1] proposed that collective motion could emerge from simple, individual-level rules: (i) attraction, the tendency to stay close to others; (ii) alignment, the inclination to move in the same direction as neighbors; and (iii) avoidance, the need to prevent collisions. Vicsek et al. [2] introduced a statistical model where individuals align their velocities with neighbors within a given proximity, incorporating stochastic noise to reflect individual variation. Couzin et al. [3] expanded these ideas by integrating spatial interaction zones, with avoidance in the innermost region, alignment in the intermediate zone, and attraction in the outermost zone.
These Attraction, Avoidance, Alignment rules (“3A”) effectively reproduce key collective behavioral phases observed in nature. We define these phases as: swarming [^1] Note, this ”swarming” phase is a different concept to swarm robotics. Here, the swarming phase is an emerging behavior that occurs when the group moves in a non-organized motion. (non-polarized cohesive motion), schooling (highly polarized motion, also known as flocking), and milling (circular motion). Models based on these principles are commonly referred to as “3A models”  [4, 5, 6, 7].
Many “3A models” rely on artificial methods to enforce rules, such as inter-individual distances [8, 9, 10, 11], global velocity measurements [12, 13], or absolute communication networks [14, 15]. However, such approaches often diverge from the mechanisms used by animals, prompting researchers to explore more biologically plausible (”bio-plausible”) models [16, 17, 18, 19]. Vision-based mechanisms have proven particularly effective, as vision is the primary sensory modality for many species [20, 21, 22, 23].
Vision-based collective motion models have previously been tested on robotic platforms [24, 25]. Moreover, some authors have previously used only vision to successfully model collective motion [26, 27, 28, 29]. Some implementations treated vision as a selective mechanism [30, 31] or do not fully adhere to the fundamental principles of the classical 3A model by not using its structure of “alignment”, “attraction” or “avoidance” rules [27, 32]. Recently, Castro et al. (2024) [33] proposed a model that relies solely on visual information while maintaining the structure of 3A models. Their model replicates collective behaviors using “early vision” cues : (i) angular position, (ii) angular size in the field of view as well as (iii) optic flow, the visual perception of the apparent angular motion of the environment caused by relative movement. Optic flow is a key navigational tool in nature, used by insects during flight [34, 35] and by fish and birds [36, 37]. Robotic implementations of optic flow have also demonstrated significant potential [38, 39, 40, 41, 42].
1.2 Contribution
@@
This study proposes a robot-in-the-loop approach to collective motion that relies exclusively on these “early vision” cues using educational spherical robots. Unlike previous models, this approach avoids reliance on global information or complex communication networks. Instead, it introduces a visual anchor to address environmental boundaries, eliminating the need for explicit wall modeling. Through numerical simulations and experiments involving 10 independently controlled robots, the method successfully reproduces diverse collective motion behaviors.
Figure 1: Modeling collective behaviors from optic flow and retinal cues [33]: a) View of the flock positions. b) Representation of flock positions after recognition at time step $t$. Each $i$-particle is described by its XY-coordinates and heading unit vector $\mathbf{e}_{i}$, with heading angle $\theta_{i}$. c) Point-of-view of particle 0, showing the relative retinal positions of each particle. The arrows represent the perceived optic flow, with the white arrow indicating the apparent velocity ($\mathbf{V}_{i,k}$). Only the azimuthal component of $\mathbf{V}_{i,k}$ is shown for retinal object $k$ as perceived by $i=0$. d) $\mathcal{V}(\phi)$: Binary panorama representation of the field of view of particle 0. e) Apparent distance $\mathcal{R}_{i}(\phi)$ for particle 0. f) Optic flow $\mathcal{O}_{i}(\phi)$ perceived by particle 1. g) Optic flow divergence $\mathcal{D}_{i}(\phi)$ perceived by particle 1.
@@
The article is organized as follows: Section 2.1 provides an overview of the challenges of a robotic implementation of 3A-models, with the example of the visual-based collective motion model proposed by Castro et al. (2024) [33] (Fig.1). Section 2.2 describes the robotic platform and the model modifications for implementing on the 10 independently controlled robots including the introduction of the visual anchor concept. Section 2.4 detail the implementation of a feature matching optic flow algorithm and visual collective motion rules. Section 2.5 describes the supplementation on original model and summarizes robotic setup, while Section 3 presents the results from simulations and robot-in-the-loop experiments. Finally, Section 4 concludes with a discussion of the results and directions for future research.
@@ 2.1 Robotic Implementation Challenges
Flocking models based on individual rules of attraction, avoidance, and alignment—commonly referred to as 3A-models—pose several challenges when implemented on robotic platforms. Two significant challenges arise due to differences between theoretical models and real-world implementations.
The first challenge is related to the abstraction of individual agents. Most 3A-models assume that individuals are self-propelled particles moving in a plane. However, many robots either deviate from the particle abstraction or lack natural constraints to planar motion. This raises the question: Which robot kinematics should be chosen? The chosen robotic platform must strike a balance between adhering to model assumptions, ensuring adequate sensing capabilities, accommodating computational constraints, and maintaining a suitable size, among other factors.
The second challenge involves transitioning from numerical simulations to physical experiments. Unlike numerical environments, experimental setups are spatially constrained. As a result, collective behaviors must be confined within the experimental area. Without addressing these constraints, implementing flocking behaviors on robots would merely involve a direct adaptation of an established model, such as the one proposed by Castro et al., (2024) [33], to the chosen robotic platform (illustrated in Fig. 1).
The following sections present our solutions to these challenges.
@@ 2.2 Robotic Setup
This section describes the robotic setup for implementing a robot-in-the-loop visual flocking model. We propose to use educational spherical robots equipped with wireless control capabilities, while the robot’s sensing capabilities are provided by an external infrastructure. The external infrastructure reconstructed each binary panorama and processed the model described in Section 2.1, controlling the locomotion of each robot independently.
@@
Figure 2: Spherical robotic platform. a) Sphero Bolt®wireless educational robots. b) Example of the robots’ movement, c) Heading control by differential rotation on the two opposite inner wheels, d) Induced nose-up pitch during forward displacement.
@@
This robot-in-the-loop solution leverages an existing educational robotic platform, the Sphero Bolt®(Fig. 2a). These compact, spherical robots can be controlled in real-time via Bluetooth®using an educational application, which allows users to set a heading, desired velocity, and LED states, as well as read onboard sensor data. Furthermore, several open-source projects extend the functionality of these robots by integrating them with programming environments such as ROS, Python, and/or MATLAB®.
@@
Each robot is equipped with a suite of sensors, including an Inertial Measurement Unit (IMU) composed of a 3-axis accelerometer, gyroscope, and magnetometer, as well as a light sensor. Additionally, the robot features four infrared (IR) proximity sensors positioned on its sides, two addressable RGB LEDs (located at the front and rear), and an 8×8 RGB LED matrix on the top.
@@
Locomotion is facilitated by a self-balancing two-wheel enclosed system, with two passive wheels (not shown in Fig.2) maintaining continuous contact with the spherical enclosure. This dual-motor configuration enables linear movement, while stationary turning is achieved via differential motion (Fig.2b-c). The locomotion process induces a pitch displacement dependent on the robot’s speed and the friction coefficient between the spherical enclosure and the surface (Fig.2d).
This pitch displacement is partially mitigated by the robot’s dual-MCU architecture: one microcontroller unit (MCU) manages the communication protocol and high-level commands, while the other is dedicated to motor control. The latter ensures precise implementation of the desired heading and speed while handling low-level sensor readings.
@@
Figure 3: Experimental Setup and external infrastructure. a) Photograph of the experimental setup. b) Sample image captured by the camera (Daheng Imaging MER2-230-168U3C camera equipped with a 4 mm C-mount lens). c) Whole-robot body image segmentation: Segmentation of the robot’s body based on the captured image. d) Recognition of individual light sources and heading: Identification of the front and rear LEDs allows calculation of the robot’s heading vector. e) Robot identification and virtualization: Assignment of a unique identifier to each robot and visualization of their positions and orientations.
@@
To address the absence of onboard cameras and the inability to integrate them into the robots, an external infrastructure made of a bird’s-eye camera was positioned above the experimental area. The camera used is an industrial RGB global shutter model, the Daheng Imaging MER2-230-168U3C, with a frame rate of 168 frames per second (FPS), a resolution of 1920×1200 pixels, and compatibility with interchangeable C-mount lenses (Table 1). A 4 mm, 1/8”, F2.0–12 non-distorting lens was selected, enabling the camera to be mounted 2.3 m above the robots while covering a 2.7×2.7 m working area—approximately 38 times the length of a single robot (Fig. 3a).
Resolution 1920x1200
Frame rate 168 FPS
@@
Non-distorting lens 4mm (focal length), 1/8” (optical format) , F2.0-12 (aperture)
Table 1: Daheng Imaging MER2-230-168U3C global shutter Camera Settings.
Each robot is assigned a unique color pattern, enabling the identification of its front and back. To maintain visual connectivity between these sections, the back half of the LED matrix is set to emit 50% bright white light. This configuration supports 16 distinct combinations for robot identification, with potential for further expansion using additional colors.
A bird’s-eye camera captures a 1920×1200 RGB image, enabling the isolation of each light source on the robots and reduces light bleeding when several robots are in close proximity (Fig. 3b). Whole-robot segmentation is performed on this image, treating the three light sources on each robot as a single entity. This allows for the calculation of the robot’s center coordinates (X,Y), which approximate its geometric center. By knowing the radius of the robots in pixels, N cropped images are extracted from these coordinates, each isolating an individual robot.
The cropped images are further segmented to recognize the front and back LEDs as well as the LED matrix. The larger area created by the two light sources facilitates accurate recognition of the robot’s front and back (Fig. 3d). Additionally, the unique color patterns from this segmentation are used to assign a distinct identifier to each robot.
Finally, the robot’s heading is calculated by tracing the vector formed between the center positions of the front and back LEDs. This computer vision processing yields the robot’s identification, X,Y-coordinates, and heading (Fig. 3e).
The next phase, following the identification and localization of each robot, involves reconstructing the binary panorama for each robot.
@@ 2.3 Anchoring the collective behavior
Figure 4: Binary Panorama Reconstruction. a) 2D virtual representation of the binary panorama: The red goal indicates the desired location of the flock. b) Mixed-reality perspective from particle 0: Displays the omniscient red goal as observed from the particle’s point of view. c) Graph representation of all $\mathcal{V}(\phi,t)$: $\mathcal{V}(\phi,t)$ represents the binary panorama view of the field of vision for particle i, with the anchor displayed as a distinct channel (red). This graph can be efficiently stored as a tuple list, representing the rising and falling edges of the binary panorama.
In the binary panorama reconstruction, we introduce a novel component to the original model [33]: a visual anchor. The visual anchor acts as an attractive force directed toward a designated location, which, in this case, is the center of the arena. The farther an individual is from this location, the stronger the anchor’s influence, causing the individual to turn and face the anchor. The anchor is described by Eq. 1, where $\mathcal{R}_{i}$ represents the apparent distance function over the binary panorama observed by particle i, and $a$ denotes the particle’s radius.
[…]
The inclusion of an anchor bounds the flock to a place and prevent infinite expansion. The anchor is an augmented reality (AR) solution that enables each robot to see the desired center point (Fig. 4a). This anchor is omnisciently visible, regardless of any obstacles between the line of sight and the anchor location (Fig. 1b). It is a new visual channel that uses the same binary representation as the visual surroundings (Fig. 1c). The binary panorama is represented in Python as an ordered tuple list of every retinal object’s rising and falling edge, with the last position always being the goal tuple. This process is performed independently for each particle without sharing information between particles.
@@ 2.4 Robotic Visual Flocking Model
Avoidance effect ($\mathcal{A}$) is here introduced in the attraction term ($\omega_{\odot}$). Our previous study [33] did not include any avoidance because most collective motion models are based on the assumption that all particles have some extra room on the z-axis, allowing them to avoid collisions with additional depth. This assumption is primarily based on how fish flocking data is gathered, where the fish can swim above or below each other in a few centimeters of water [43]. However, for any robotic implementation, obstacle or collision avoidance is of critical importance. Fortunately, avoidance and attraction have similar characteristics, being proportional instantaneous responses to the optical size and location of a retinal object. The key difference is in the direction of the response: while attraction is positive, avoidance is negative. Therefore, avoidance is implemented by providing a new piecewise constant function ($\mathcal{A}$), where attraction is inverted in situations where a particle should avoid a retinal object. This new function $\mathcal{A}$ depends on the width of the retinal object (optical size), specifically the difference between its rising and falling edges ($\Delta\phi$). Note that the apparent distance of a retinal object ($\mathcal{R}(\phi,t)$) inversely depends on this optical size $\Delta\phi$. We choose $\mathcal{A}(\Delta\phi)=-5$ for $\Delta\phi<\pi/10$, which is equivalent to the retinal object being closer than three particle radii ($\mathcal{R}<3a$). In all other cases, $\mathcal{A}(\Delta\phi)=1$. The blind spot remains unchanged, giving more weight to the retinal object in front of the particle, regardless of this new function.
@@
$\frac{\mathbf{e}_{i}(t)\times\mathbf{V}_{ik}(t)}{\mathcal{R}_{i}(t)}\cdot\hat{\mathbf{z}}=-\mathcal{D}_{i}(\phi,t)\sin(\phi,t)+\overline{\mathcal{O}}_{i}(\phi,t)\cos(\phi,t).$ (2f)
Where $\mathbf{x}_{i}(t)$ is the XY-coordinate of particle i at time t, $\mathbf{e}_{i}(t)$ is the heading vector of particle i at time t, $k_{\odot},\,k_{\parallel},\,k_{c},\,k_{\eta}$ are the gains of the attraction, parallel, anchor and noise respectively while $\omega_{\odot}(t),\,\omega_{\parallel}(t),\,\omega_{c}(t),\,\eta$ are the attraction, parallel, anchor and noise terms at timet. $\hat{\mathbf{z}}$ is the unitary vector that is orthogonal to the ($\hat{\mathbf{x}}$, $\hat{\mathbf{y}}$)-plane of movement of the robots. $\phi$ is the retinal angle in the field of view, is considered from the point of view of each particle and counter-clockwise is considered positive. $\mathcal{R}_{i}(\phi,t)$ is the apparent distance function at a retinal angle $\phi$ at time t for individual i. $b_{\epsilon}(\phi,t)$ is the blindspot function ($1+\epsilon\cos\phi$). $\mathbf{V}_{ik}(t)$ is the apparent velocity that an individual i perceives of a retinal object k. Finally $\tau$ is the delay between perception and action.
The robotic platform introduces several critical processing factors – Bluetooth, image processing, internal Spheros control command, … – that affect the overall dynamical response of the robots and its collective behavior as a whole (see Supplementary Information for more details). These extra factors results in a time delay $\tau$ with a median of 600ms. The overall functional diagram of the collective motion model is presented in Fig. 5. This model is executed independently and simultaneously for all robots; it does not involve communication between robots, sharing information, and only uses visual data. The same block could be utilized on a robotic platform using LiDAR, onboard cameras, touch sensors, ultrasound, or any other sensor that provides a binary visual panorama as output. All the information used by this model is self-referred and self-contained; there is no global reference frame, coordination, or knowledge of the flock or the environment.